1.




Given the following demand and supply function of milk in a market.




Q_{d} = 28 - 4P




+ P respectively.




Q_{s} = 18




i.




and




Determine the equilibrium price and quantity of milk?




ii.




If government fixes price at GHC 1.00, find the quantity demanded and supplied of milk and comment on the situation market. in the




What is the full economic price that consumers would end up paying as a result of (ii) above?

1
Expert's answer
2021-12-16T18:14:50-0500

We are given that,

Qd = 28 - 4P

Qs = 18 + P 

(i.) Equilibrium price and Quantity

At the equilibrium level, 

Qd = Qs 

Now, plugging Qs and Qd in above,

28 - 4P = 18 + P

10 = 5P

P = 2

Equilibrium price is GHC 2

Putting value of P in either demand or supply function to get equilibrium quantity.

Q = 18 + 2 = 20

Equilibrium quantity is 20.


(ii.) Govt has set the prices at GHC 1

Now we will find the quantity demanded and supplied at this level

 

Qd = 28 - 4P = 28 - 4(1) = 28 - 4 = 24

Qs = 18 + P = 18 + 1 = 19

whereas quantity demanded and supplied is at 20 units.

We can see that due to the reduction in prices,

Quantity demanded has increased from 20 to 24 units and the quantity supplied has decreased from 20 to 19 units.


As of now, only 19 units are supplied by producers,

So, the consumers can demand only 19 units

Qd will be 19 units. Plugging it in demand equation to find out the end up price consumer will eventually pay.

Qd = 28 - 4P = 19

4P = 9

P = 9/4 = 2.25